# Efficient and Effective Adaptation of Multimodal Foundation Models in Sequential Recommendation

### v1

@@ VI-C Performance on Public Benchmark (RQ3)
To further evaluate IISAN-Versa, we compare it with the public benchmark MicroLens, as detailed in Table V. This public benchmark [24] includes four main types of baselines: sequential recommendation (IDRec (SR)), collaborative filtering (IDRec (CF)), and ID-based recommendation with features generated from frozen pre-trained video encoder (VIDRec (Frozen Encoder)) or end-to-end (E2E) learning encoders (VideoRec (E2E learning)). The previous state-of-the-art approach, GRU4RecV [24], utilizes GRU4Rec as the sequential encoder with a tunable video encoder VideoMAE to encode video content in an E2E manner directly. To make sure a fair comparison we directly adopt the data from the original benchmark [24]. We also perform a comprehensive comparison by including a multimodal full fine-tuning of the multimodal VideoRec framework, as applied in [18].
@@
As shown in Table V, both IISAN-Versa approaches outperform existing approaches on this dataset. Notably, IISAN-Versa can significantly reduce training costs compared to VideoRec, which requires over four times more GPU resources than even full fine-tuning of the multimodal paradigm. This is due to the high computational cost of video encoders that process multiple frames during training. Our approach, with its caching strategy, results in IISAN-Versa using vastly smaller GPU memory. To further validate IISAN-Versa on MicroLens, we also incorporate the standard full fine-tuning method, which underperforms the IISAN-Versa. Within the IISAN-Versa paradigm, IISAN-VA demonstrates better performance compared to IISAN-VS, which is consistent with our previous findings.
(Answer to RQ3) IISAN-Versa achieves state-of-the-art performance on the public benchmark MicroLens dataset.
@@ VI-D Efficiency Analysis (RQ4)
TABLE VI: Ablation study for IISAN-Versa and IISAN-VA on Scientific Dataset. IISAN-Versa mainly contains four key components: LayerDrop, Modality Gate, and Intra- and Inter-modal towers. ”Frozen Backbone” refers to keeping the modality backbone frozen, without applying any adaptation methods.
@@
[…]
TABLE VII: LayerDrop in symmetrical IISAN-Versa on Scientific Dataset. The number of blocks in the Method column represents that keep this number of blocks and drop the others.
@@
TABLE VIII: Ablation Study for blocks of asymmetrical IISAN-Versa. n(even_all) refers to selecting n blocks evenly distributed across all blocks. ”Top” and ”down” indicate that the blocks are drawn from either the upper half or the lower half, respectively.
Combination Configuration HR@10 NDCG@10
@@
12+6 12(even_all)+6(even_all) 7.08 4.44
12(even_all)+6(top) 6.80 4.19
12(even_all)+6(down) 7.15 4.29
@@
6+6 6(even_all)+6(even_all) 7.29 4.41
6(even_all)+6(down) 7.21 4.37
6(even_all)+6(top) 6.80 4.19
6(even_top12)+6(even_all) 6.96 4.24
6(even_top12)+6(top) 6.95 4.27
6(even_top12)+6(down) 6.97 4.25
6(even_down12)+6(even_all) 7.29 4.44
6(even_down12)+6(top) 6.87 4.26
6(even_down12)+6(down) 7.35 4.51
@@
TABLE IX: Performance on Multimodal Text.
@@ VI-E Ablation Studies (RQ5)
In this section, we first demonstrate the importance of the module’s layer dropping for both symmetrical and asymmetrical IISAN-Versa. The BERT-large is applied as the text encoder for IISAN-VA’s ablation study. We extensively explore the blocks required for asymmetrical IISAN-Versa. Specifically, we experiment with the combinations of blocks, ranging from {24+12, 12+12, 12+6, 6+6}, with the former value in each combination indicating the blocks for 24-layer BERT-large and the latter representing the blocks for 12-layer ViT-base,
@@
(1) Modality Selection. Table VI highlights that using separate intra-modal SAN (Line 4) or inter-modal SAN (Line 5) clearly reduces recommendation effectiveness. However, compared to the pre-trained backbones (Line 3) with frozen layers (where only the sequential encoder is trained), both intra-modal SAN and inter-modal SAN can significantly improve performance, performing best when used simultaneously. This reflects that intra- and inter-modal SAN can enhance feature adaptation within each modality and improve inter-modal feature interaction, such argument holds for both IISAN-VS and IISAN-VA.
@@
(2) Gated Fusion. To deepen our understanding of the differences in information usage between textual and visual modalities, we analyzed the gate weights at the optimal checkpoint for both symmetrical and asymmetrical IISAN-Versa. These gate values, ranging from 0 to 1, imply that values above 0.5 denote a stronger focus on a specific modality. The data indicate that in multimodal recommendations, the weights assigned to visual modality gates consistently range from 0.2 to 0.4 in all structures. This trend suggests a predominant reliance on textual modality within our multimodal approach.
@@
(3) LayerDrop Strategies for IISAN-Versa. IISAN-VS’s layers exhibit a degree of redundancy, allowing LayerDrop to enhance performance effectively. First, when we remove LayerDrop, both HR@10 and NDCG@10 decrease by 0.1%, demonstrating LayerDrop’s effectiveness. Second, we study the effect of different LayerDrop strategies, as described in Table VII. We adopt the even-numbered $(2,4,6,\ldots,12)$ transformer blocks (6 blocks), skipping the odd-numbered layers, which achieves the best performance-efficiency balance. In terms of IISAN-VA, the performance drop is even more significant when LayerDrop is completely removed, indicating that the redundancy is more prominent than in IISAN-VS. To explore how many blocks should be used for IISAN-VA, we perform extensive experiments on the combination of the layers and compare their mean values. Due to the complexity of the combination, we study the even selection for layer group of {top, down, all} for the 6+6 combination. When the numbers are asymmetrical such as “24+12” and “12+6”, we adopt the top hidden states to align with the vision tower for Inter-SAN. The intra-modality stays the same. Extensive experiments show that the “6+6” structure has the best overall mean performance among all structures. This further supports the point that IISAN-VA has more layer redundancy compared to IISAN-VS, due to its greater number of incorporated layers. [^8] Exhaustively testing all layer combinations, even for smaller models like IISAN-VA (BERT-large + ViT-base), would require about 68 billion evaluations. Instead, we aim to find a general strategy for larger models. We adopt the 6(even_all) + 6(even_all) method as the default for IISAN-VA, which, while not the best, performs above average among 6+6 combinations in Table VIII, offering convenience and generality for asymmetrical structures.
@@ VI-F Performance on Multimodal Text (RQ6)
The results, shown in Table IX, demonstrate that IISAN-Versa effectively adapts to multimodal text by outperforming both title-based and caption-based approaches. This superior performance over these three baselines highlights IISAN-Versa’s versatility. Additionally, the combination of raw text—title and video caption—proves to be the most informative. By learning from this combination of these texts, we achieve the best overall performance in multimodal text scenarios. However, it is important to note that no combination of multimodal text outperforms the use of titles with raw cover images [^9] Note that involving raw video and audio are beyond the scope of this paper., indicating that raw modality remains crucial in the multimodal representation of sequential recommendations.
@@ VII Conclusion and Future Work
This paper introduces the new IISAN-Versa, an extension of the IISAN framework, which offers a versatile and efficient architecture for multimodal sequential recommendation tasks. IISAN-Versa overcomes the previous limitation of being applicable only to large asymmetrical MFMs, demonstrating a clear scaling effect where larger text models consistently enhance recommendation performance. Additionally, IISAN-Versa proves effective in handling diverse multimodal text scenarios, significantly broadening its applicability. Future research should aim to overcome current limitations by incorporating additional modalities such as audio, raw video, or graph information. Furthermore, exploring the potential of adapting the IISAN-Versa framework to other tasks, such as Visual Question Answering (VQA) and video retrieval, presents an intriguing avenue for future work. Another promising research direction involves optimizing training efficiency by reducing the number of trainable parameters, particularly in large text encoders like Llama-70B, given their extensive embedding dimensions.

### v2

@@ VI-C Performance on Public Benchmark (RQ3)
To further evaluate IISAN-Versa, we compare it with the public benchmark MicroLens, as detailed in Table VII. This public benchmark [18] includes four main types of baselines: sequential recommendation (IDRec (SR)), collaborative filtering (IDRec (CF)), and ID-based recommendation with features generated from frozen pre-trained video encoder (VIDRec (Frozen Encoder)) or end-to-end (E2E) learning encoders (VideoRec (E2E learning)). The previous state-of-the-art approach, GRU4RecV [18], utilizes GRU4Rec as the sequential encoder with a tunable video encoder VideoMAE to encode video content in an E2E manner directly. To make sure a fair comparison we directly adopt the data from the original benchmark [18]. We also perform a comprehensive comparison by including a multimodal full fine-tuning of the multimodal VideoRec framework, as applied in [14].
@@
Figure 5 and Table 5 presents a comparison between our proposed model and the widely adopted ID-based baseline SASRec, focusing on both effectiveness and computational efficiency. Although SASRec demonstrates strong performance in sequential recommendation, our model consistently achieves higher accuracy across all metrics while requiring less GPU memory. While multimodal processing inherently introduces additional training time, the combination of reduced memory consumption and improved accuracy underscores the practical potential of our approach. These results affirm the real-world viability of our method, particularly in deployment scenarios where both performance and GPU memory are essential.
@@
As shown in Table VII, both IISAN-Versa approaches outperform existing approaches on this dataset. Notably, IISAN-Versa can significantly reduce training costs compared to VideoRec, which requires over four times more GPU resources than even full fine-tuning of the multimodal paradigm. This is due to the high computational cost of video encoders that process multiple frames during training. Our approach, with its caching strategy, results in IISAN-Versa using vastly smaller GPU memory. To further validate IISAN-Versa on MicroLens, we also incorporate the standard full fine-tuning method, which underperforms the IISAN-Versa. Within the IISAN-Versa paradigm, IISAN-VA demonstrates better performance compared to IISAN-VS, which is consistent with our previous findings.
(Answer to RQ3) IISAN-Versa surpasses state-of-the-art performance on the public benchmark MicroLens dataset.
@@ VI-D Efficiency Analysis (RQ4)
TABLE IX: Ablation study for IISAN-Versa and IISAN-VA on Scientific Dataset. IISAN-Versa mainly contains four key components: LayerDrop, Modality Gate, and Intra- and Inter-modal towers. Frozen Backbone” refers to keeping the modality backbone frozen, without applying any adaptation methods.
@@
IISAN-Versa 6.83 4.14 7.29 4.41
Figure 9: Visualization of inter-modality gates in IISAN-Versa. Darker colors indicate higher gate weights.
TABLE X: LayerDrop in symmetrical IISAN-Versa on Scientific Dataset. The number of blocks in the Method column represents that keep this number of blocks and drop the others.
@@
TABLE XI: Ablation Study for blocks of asymmetrical IISAN-Versa. n(uniform_all) refers to selecting n blocks evenly distributed across all blocks. Top” and down” indicate that the blocks are drawn from either the upper half or the lower half, respectively.
Method Configuration HR@10 NDCG@10
@@
12+6 12(uniform_all)+6(uniform_all) 7.08 4.44
12(uniform_all)+6(top) 6.80 4.19
12(uniform_all)+6(down) 7.15 4.29
@@
6+6 6(uniform_all)+6(uniform_all) 7.29 4.41
6(uniform_all)+6(down) 7.21 4.37
6(uniform_all)+6(top) 6.80 4.19
6(uniform_top12)+6(uniform_all) 6.96 4.24
6(uniform_top12)+6(top) 6.95 4.27
6(uniform_top12)+6(down) 6.97 4.25
6(uniform_down12)+6(uniform_all) 7.29 4.44
6(uniform_down12)+6(top) 6.87 4.26
6(uniform_down12)+6(down) 7.35 4.51
@@
TABLE XII: Performance on Multimodal Text.
@@ VI-E Ablation Studies (RQ5)
In this section, we first demonstrate the importance of the module’s layer dropping for both symmetrical and asymmetrical IISAN-Versa. The BERT-large is applied as the text encoder for IISAN-VA’s ablation study. We extensively explore the blocks required for asymmetrical IISAN-Versa. Specifically, we experiment with the combinations of blocks, ranging from {24+12, 12+12, 12+6, 6+6}, with the former value in each combination indicating the blocks for 24-layer BERT-large and the latter representing the blocks for 12-layer ViT-base.
@@
(1) Modality Selection. Table IX highlights that using separate intra-modal SAN (Line 4) or inter-modal SAN (Line 5) clearly reduces recommendation effectiveness. However, compared to the pre-trained backbones (Line 3) with frozen layers (where only the sequential encoder is trained), both intra-modal SAN and inter-modal SAN can significantly improve performance, performing best when used simultaneously. This reflects that intra- and inter-modal SAN can enhance feature adaptation within each modality and improve inter-modal feature interaction, such argument holds for both IISAN-VS and IISAN-VA.
@@
(2) Gated Fusion. To deepen our understanding of the differences in information usage between textual and visual modalities, we analyzed the gate weights at the optimal checkpoint for both symmetrical and asymmetrical IISAN-Versa. We visualize the gate values in Figure 9. Since text gate values are consistently above 0.5, all models primarily rely on text. Visual gate values stay between 0.26 and 0.42, with IISAN-VA depending on text even more than IISAN-VS. This trend suggests a predominant reliance on textual modality within our multimodal approach.
@@
(3) LayerDrop Strategies for IISAN-Versa. IISAN-VS’s layers exhibit a degree of redundancy, allowing LayerDrop to enhance performance effectively. First, when we remove LayerDrop, both HR@10 and NDCG@10 decrease by 0.1%, demonstrating LayerDrop’s effectiveness. Second, we study the effect of different LayerDrop strategies, as described in Table X. We adopt the even-numbered $(2,4,6,\ldots,12)$ transformer blocks (6 blocks), skipping the odd-numbered layers, which achieves the best performance-efficiency balance. In terms of IISAN-VA, the performance drop is even more significant when LayerDrop is completely removed, indicating that the redundancy is more prominent than in IISAN-VS. To explore how many blocks should be used for IISAN-VA, we perform extensive experiments on the combination of the layers and compare their mean values (see Table XI). Due to the complexity of the combination, we study the even selection for layer group of {top, down, all} for the 6+6 combination. When the numbers are asymmetrical such as “24+12” and “12+6”, we adopt the top hidden states to align with the vision tower for Inter-SAN. The intra-modality stays the same. Extensive experiments show that the “6+6” structure has the best overall mean performance among all structures. This further supports the point that IISAN-VA has more layer redundancy compared to IISAN-VS, due to its greater number of incorporated layers. [^8] Exhaustively testing all layer combinations, even for smaller models like IISAN-VA (BERT-large + ViT-base), would require about 68 billion evaluations. Instead, we aim to find a general strategy for larger models. We adopt the 6(uniform_all) + 6(uniform_all) method as the default for IISAN-VA, which, while not the best, performs above average among 6+6 combinations in Table XI, offering convenience and generality for asymmetrical structures.
@@ VI-F Performance on Multimodal Text (RQ6)
The results, shown in Table XII, demonstrate that IISAN-Versa effectively adapts to multimodal text by outperforming both title-based and caption-based approaches. This superior performance over these three baselines highlights IISAN-Versa’s versatility. Additionally, the combination of raw text—title and video caption—proves to be the most informative. By learning from this combination of these texts, we achieve the best overall performance in multimodal text scenarios. However, it is important to note that no combination of multimodal text outperforms the use of titles with raw cover images [^9] Note that involving raw video and audio are beyond the scope of this paper., indicating that raw modality remains crucial in the multimodal representation of sequential recommendations.
@@
VI-G Hyper-parameter Sensitivity Analysis (RQ7)
In this section, we address RQ7 by conducting a sensitivity analysis on two key hyperparameters—learning rate and embedding dimension—following the setup in [10], to provide a more comprehensive empirical evaluation of PEFT-based modality recommendation.
As shown in Figure 10, both IISAN-Versa variants demonstrate robust performance across a range of hyperparameter settings. IISAN-VS maintains stable performance regardless of the learning rate or embedding dimension. IISAN-VA also remains stable overall, but its performance slightly degrades when either the embedding dimension or the learning rate is set too high. The best performance is achieved when both hyperparameters are set to relatively small values; notably, smaller embedding dimensions can enhance efficiency.
(Answer to RQ7) Both IISAN-VS and IISAN-VA exhibit robustness to hyperparameter variations. However, IISAN-VA requires more conservative tuning, with smaller embedding dimensions and moderate learning rates yielding improved performance and efficiency.
Figure 10: Learning rate and Embedding Dimension of IISAN-Versa. The first row represents the performance of IISAN-VS on the Scientific dataset whereas the second row is for IISAN-VA.
@@ VII Conclusion and Future Work
This paper introduces the new IISAN-Versa, an extension of the IISAN framework, which offers a versatile and efficient architecture for multimodal sequential recommendation tasks. IISAN-Versa overcomes the previous limitation of being applicable only to large asymmetrical MFMs, demonstrating a clear scaling effect where larger text models consistently enhance recommendation performance. Additionally, IISAN-Versa proves effective in handling diverse multimodal text scenarios, significantly broadening its applicability.
Future research should aim to overcome current limitations by incorporating additional modalities such as audio, raw video, or graph information. Furthermore, exploring the potential of adapting the IISAN-Versa framework to tasks such as Visual Question Answering (VQA) and video retrieval is a promising direction, as these tasks commonly adopt dual-tower architectures where one tower encodes the visual modality and the other the textual input. IISAN-Versa naturally fits into this paradigm, with its modality-specific intra-SANs and Inter-SAN facilitating effective interaction between different modalities. Extending to additional modalities like audio or video can be easily achieved by adding more intra-SANs for each modality and including them in the Inter-SAN computation, enabling flexible multimodal fusion while maintaining adaptation efficiency. Another promising research direction involves optimizing training efficiency by reducing the number of trainable parameters, particularly in large text encoders like Llama-70B, given their extensive embedding dimensions.